How many cups of pure granola must be added to 16 cups of granola that is 15% raisins so that the percent of raisins in the resulting mixture will be 10%?

Answer by stanbon(75887)

How many cups of pure granola must be added to 16 cups of granola that is 15% raisins so that the percent of raisins in the resulting mixture will be 10%?
---------------------------
Equations:
raisins + raisins = raisins
---
0*x + 0.15*16 = 0.10(x+16)
---
Multiply thru by 100 to get:
0 + 15*16 = 10x + 10*16
10x = 5*16
x = 8 cups (amt. of pure granola needed for the mixture)
=========================
Cheers,
Stan H.